# DocLayLLM: An Efficient and Effective Multi-modal Extension of Large Language Models for Text-rich Document Understanding

### v2

@@ 4.4 Comparisons with SOTA OCR-free methods
We have also compared our method with SOTA OCR-free TDU models. Table 4 presents the precision of our method and some advent OCR-free TDU methods across different benchmarks. As shown in the table, even though in the circumstance that these OCR-free models typically use the corresponding training sets of these benchmarks during training, while our method infers in a purely zero-shot manner, our DocLayLLM still outperforms these models, particularly in the KIE task, which heavily relies on layout information. This clearly demonstrates the superiority of our design which provides LLM with OCR information.
@@
We additionally test DocLayLLM on a broader set of benchmarks from DocOwl 1.5 Ye et al. (2023a). To further assess the performance of DocLayLLM on seen datasets, we incorporate the training sets of these downstream benchmarks used in DocOwl 1.5 during the SFT process. It is worth noting that OCR-free TDU methods typically include these datasets during training. To maintain the efficiency advantage of our method, we ensure that the SFT is conducted with the same number of steps as the standard DocLayLLM. The results in Table 5 show that with the inclusion of downstream task data during SFT, our method outperforms the SOTA OCR-free TDU models in most datasets. Moreover, our overall performance in the Document-oriented VQA, KIE, and Table Understanding sub-tasks surpasses the OCR-free TDU methods using limited training resources, showcasing the tremendous potential of OCR-dependent methods and further reinforcing the reliability of our method.
@@
Overall, even with fewer training resources, our DocLayLLM shows strong zero-shot inference performance, outperforming many OCR-free TDU methods without prior exposure to downstream datasets. Additionally, when incorporating downstream datasets during SFT like other OCR-free approaches, DocLayLLM surpasses them, highlighting its potential in solving TDU tasks by integrating OCR information into the LLM. For the visualization of qualitative examples, please refer to Appendix A.2.
@@
w/o CoT Pre-training 76.00 350.18 213.09 24.54 25.49 25.02
w/o CoT Annealing 77.40 353.08 215.24 34.23 26.21 30.22
w/o Patch Features 77.32 356.76 217.04 33.75 25.97 29.86
Random $VE$ $LE$ 77.42 355.54 216.48 31.31 25.76 28.54
DocLayLLM 77.43 358.68 218.06 35.41 26.25 30.83
Table 6: Ablation study on various design components of our DocLayLLM.
4.5 Ablation Study
We have also conducted additional experiments to explore the effectiveness of our proposed method. To conserve computational resources, we perform pre-training at just 25% of the standard scale. Our primary focus is to validate the structure of our design and the impact of the proposed CoT Pre-training and CoT Annealing techniques. The main results are presented in Table 6, where w/o CoT Pre-training and w/o CoT Annealing mean pre-training without CoT and supervised fine-tuning only with CoT data. w/o Patch features denotes the absence of patch features and the corresponding pre-training task, and Random $VE$ and $PE$ refers to the use of randomly initialized visual embedder and layout embedder.
The experimental results indicate that omitting CoT Pre-training leads to an obvious performance drop in both the Document-oriented VQA and KIE tasks, while not using CoT Annealing primarily affects the performance of the Document-oriented VQA task. Besides, excluding patch features or using randomly initialized $VE$ and $LE$ partially compromise the proposed extension, resulting in performance degradation in both the Document-oriented VQA and KIE tasks. Overall, the combined impact of these components ensures the efficiency and effectiveness of our method.
5 Conclusion
@@
We propose DocLayLLM, an efficient and effective multi-modal extension of large language models for text-rich document understanding. By inserting document patch tokens and 2D positional tokens into the LLM through the simple linear transformation, it achieves an efficient multi-modal extension, enabling the LLM to perceive and process textual, visual, and layout information from the OCR results of the document. Additionally, we introduce CoT Pre-training and CoT Annealing techniques centered around CoT, further enhancing the LLM’s ability to handle OCR-derived information, thereby strengthening its understanding of text-rich documents. Despite using less data and fewer training resources, our method surpasses existing OCR-dependent TDU models, demonstrating strong zero-shot generalization capability. Besides, it also exceeds the performance of OCR-free TDU methods in the same multi-downstream-task fine-tuning setting, showcasing exceptional performance and further underscoring its efficiency and effectiveness.
For the future work, we aim to further unlock the power of DocLayLLM and explore involving more types of documents during our model learning, to develop a more versatile, robust, and powerful document understanding model.
Appendix A Appendix
A.1 More Implement Details
@@
We implement our DocLayLLM using Llama2-7B-Chat (Touvron et al., 2023) and Llama3-8B-Instruct (Llama Team, 2024). The hyper-parameters for both pre-training and supervised fine-tuning are detailed in Table 7. As shown in the table, our DocLayLLM demonstrates efficiency, requiring fewer training resources while maintaining high performance. This underscores the method’s capability to deliver robust results without the need for extensive computational power, making it a resource-efficient solution for text-rich document understanding tasks.
@@
Table 7: Hyper-parameters of DocLayLLM.
A.2 Qualitative Examples
We also provide some qualitative examples to visualize the performance of our DocLayLLM compared with other methods. Figure 2 shows the visualizations of our DocLayLLM, LayTextLLM (Lu et al., 2024), and DocOwl 1.5 (Hu et al., 2024a) on KIE benchmarks. Whether compared to OCR-free or OCR-dependent methods, it is evident that our DocLayLLM excels at handling documents with complex structures. Additionally, it more accurately locates answers and generates complete responses to questions.
We further present more visualizations across a wider range of benchmarks. In these comparisons, our DocLayLLM, using a portion of mixed downstream task data during the supervised fine-tuning process, is evaluated against the SOTA OCR-free method, which undergoes a similar fine-tuning process. As illustrated in Figure 3, the results highlight DocLayLLM’s superior capability in processing tables and its reduced hallucinations compared to DocOwl 1.5. This demonstrates the significant potential of our OCR-dependent approach in addressing text-rich document understanding tasks and underscores the rationale behind our proposed extension.
A.3 Preliminary Comparisons with MLLMs
Typically, many OCR-free TDU models are built upon an MLLM. With the rapid advancement of MLLMs, they have also shown impressive performance in text-rich document understanding. Therefore, we conduct a preliminary comparison between our DocLayLLM and some of the most powerful MLLMs available today, including the open-source InternVL2 (OpenGVLab, 2024) and the closed-source GPT-4o (OpenAI, 2024). These models are often evaluated on OCRBench (Liu et al., 2023b) to test their OCR capabilities. The OCRBench includes the KIE sub-task, consisting of 66 samples from FUNSD, 67 from SROIE, and 67 from POIE, amassing a total of 200 test samples for document question-answering. These are designed to evaluate the model’s key information extraction capabilities on text-rich documents. Thus, we compare our DocLayLLM against these models on the KIE sub-task in OCRBench.
Figure 2: Qualitative comparisons with OCR-free and OCR-dependent methods on KIE benchmarks, including FUNSD (Jaume et al., 2019), SROIE (Huang et al., 2019), and POIE (Kuang et al., 2023).
@@ A.3 Preliminary Comparisons with MLLMs
Figure 3: Qualitative comparisons with DocOwl 1.5 across various benchmarks. The document-oriented VQA tasks include DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022), and VisualMRC (Tanaka et al., 2021); the KIE task includes DeepForm (Svetlichnaya, 2020); and the Table Understanding task includes WTQ (Pasupat and Liang, 2015). indicates that the model uses the training set corresponding to the evaluation set during training.
As shown in Table 8, even when compared with the most powerful MLLMs, our DocLayLLM, using significantly fewer training resources and a smaller parameter size, performs on par with or even outperforms them in TDU tasks. This highlights the robust performance of our proposed efficient and effective extension, further demonstrating the immense potential of OCR-dependent solutions in addressing text-rich document understanding challenges. It also underscores the significant promise of combining traditional OCR techniques with advanced LLMs.
Method KIE in OCRBench
FUNSD SROIE POIE In Total
InternVL2-8B (OpenGVLab, 2024) 52 49 59 160
GPT-4o (OpenAI, 2024) 53 55 61 169
DocLayLLMLlama3 (Ours) 62 58 58 178
Table 8: Comparisons with SOTA MLLMs on the KIE sub-tasks within OCRBench Liu et al. (2023b). The metric reflects the number of samples correctly answered.

### v3

@@ 4.4 Comparisons with SOTA OCR-free methods
We have also compared our method with SOTA OCR-free TDU models. Table 3 presents the precision of our method and some advent OCR-free TDU methods across various benchmarks. Even though in the circumstance that OCR-free models use some of the training sets of these benchmarks during training, while our method infers in a purely zero-shot manner, our DocLayLLM still outperforms these models, particularly in KIE task, which heavily relies on layout information. This clearly demonstrates the superiority of our design to provide LLM with OCR information.
@@
We additionally test DocLayLLM on a broader set of benchmarks from DocOwl 1.5 [17]. To further assess the performance of DocLayLLM on seen datasets, we incorporate the training sets of these benchmarks used in DocOwl 1.5 during the SFT process as other OCR-free TDU methods do during training. To maintain the efficiency advantage of our method, we ensure that the SFT is conducted with the same number of steps as the standard DocLayLLM. The results in Table 4 show that with the inclusion of downstream task data during SFT, our method outperforms the SOTA OCR-free TDU models in most datasets. Moreover, our overall performance in the Document-oriented VQA, KIE, and Table Understanding sub-tasks surpasses the OCR-free TDU methods using limited training resources, showcasing the tremendous potential of OCR-dependent methods and further reinforcing the reliability of our method.
@@
Overall, our DocLayLLM consistently shows strong zero-shot inference ability, outperforming many OCR-free TDU methods without prior exposure to downstream data. Additionally, when incorporating downstream datasets during SFT like other OCR-free approaches, DocLayLLM surpasses them, highlighting its potential in solving TDU tasks by integrating OCR information into the LLM.
Figure 3: Qualitative comparisons of DocLayLLM with other OCR-dependent and OCR-free methods. Zoom in for better view.
4.5 Qualitative Results
Figure 3 shows some qualitative examples to visualize the performance on KIE tasks of our DocLayLLM compared with LayTextLLM and DocOwl 1.5. The example from SROIE clearly shows that DocLayLLM has improved document comprehension and answer accuracy. And the example from POIE demonstrates our advantage in handling table-formatted documents to give an accurate response. Furthermore, as seen in the FUNSD examples, DocLayLLM excels at extracting answers spanning multiple text lines, outperforming both OCR-dependent and OCR-free methods. For more visualization of qualitative examples, please refer to the appendix.
4.6 Ablation Study
We have also conducted additional experiments to explore the effectiveness of our proposed method. To conserve computational resources, we perform pre-training at 25% of the standard scale (including the baseline) but full-scale SFT. We focus on the impact of two main factors: (a) the role of CoT pre-training and CoT annealing, and (b) the influence of different modality configurations.
Experimental results in Table 5(a) indicate that omitting CoT Pre-training leads to an obvious performance drop in both the Document-oriented VQA and KIE tasks, while not using CoT Annealing primarily affects VQA performance. Besides, using randomly initialized $VE$ and $LE$ partially compromises the proposed extension, resulting in performance degradation in both the VQA and KIE tasks.
Additionally, performances in Table 5(b) show the efficacy of our DocLayLLM to incorporate different modalities for TDU tasks. It is obvious that using only text from the document is insufficient for TDU, while adding layout information in textual form provides some gain on VQA tasks but leads to over-length input (see appendix for analysis), greatly slowing the inference. By our design to integrate layout features via an LE, we achieve notable improvements on these TDU tasks with shorter inputs. And adding visual features further boosts the performance of our method.
Overall, the combined impact of these designs ensures the efficiency and effectiveness of our method.
5 Conclusion
@@
DocLayLLMLlama3 77.43 316.27 196.85 35.41 26.25 30.83
    w/o CoT Pre-training 76.00 295.49 185.75 24.54 25.49 25.02
    w/o CoT Annealing 77.40 310.17 193.79 34.23 26.21 30.22
    w/o pt VE & LE 77.42 308.49 192.96 31.31 25.76 28.54
(a) Impact of different settings: w/o CoT Pre-training and w/o CoT Annealing indicate pre-training without CoT and supervised fine-tuning with only CoT data, while w/o pt VE and PE uses randomly initialized visual and layout embedders instead of pre-trained ones.
Modality Document-oriented VQA KIE
T L V DocVQA VisualMRC Avg. DeepForm KLC Avg.
73.57 280.27 176.93 15.76 24.09 19.93
(I) 75.72 306.51 191.12 14.13 24.23 19.18
(II) 77.32 310.75 194.03 33.75 25.97 29.86
(II) 77.43 316.27 196.85 35.41 26.25 30.83
(b) Impacts of different modality configurations. T, L, and V represent textual, layout, and visual modalities, respectively. (I) indicates inputting OCR coordinates as text following He et al. [14], while (II) embeds coordinates with $LE$ in our approach.
Table 5: Ablation study on various designs of our DocLayLLM.
@@
We propose DocLayLLM, an efficient multi-modal extension of large language models for text-rich document understanding. By streamlined integration of document patch tokens and OCR positional tokens into the LLM inputs, it achieves a simple yet efficient multi-modal extension of LLM to enable it to have the capability to perceive and process textual, visual, and layout information from the OCR results of the document. Additionally, we introduce CoT Pre-training and CoT Annealing techniques centered around CoT, offering a new way to generate CoT data and further enhancing the model’s understanding of text-rich documents. Despite requiring fewer training resources, our method surpasses existing OCR-dependent TDU models, demonstrating strong zero-shot generalization capability. Furthermore, it also exceeds the performance of SOTA OCR-free TDU methods under comparable training conditions, showcasing remarkable performance across various document understanding tasks. Our work significantly advances efficient MLLM-based document understanding while maintaining high-performance standards.
Acknowledgement
This research is supported in part by National Natural Science Foundation of China (Grant No.: 62476093, 62441604).
Supplementary Material
A Pre-traing tasks
We adopt pre-training tasks as shown in Table A. These tasks facilitate the alignment of layout and visual features with the LLM’s feature space while enhancing the LLM’s understanding of document content.
Task Description
Document Description Provide a brief overview of the document.
Layout Analysis Determine the layout type (e.g., Title, Author, Paragraph, etc) of a giving area or locate specific layout elements.
Mask Language Model Restore masked words in the provided OCR text.
Table A: The training tasks during the pre-training stage.
B More Implementation Details
@@
We implement our DocLayLLM using Llama2-7B-Chat [64] and Llama3-8B-Instruct [42]. The hyper-parameters for both pre-training and supervised fine-tuning are detailed in Table B. As shown in the table, our DocLayLLM demonstrates efficiency, requiring fewer training resources while maintaining high performance. This underscores the method’s capability to deliver robust results without the need for extensive computational power, making it a resource-efficient solution for text-rich document understanding tasks.
@@
Table B: Hyper-parameters of DocLayLLM.
C More Qualitative Examples
We also provide additional qualitative examples of our DocLayLLM. As shown in the comparison between DocLayLLM and the SOTA OCR-free method DocOwl 1.5 [17] in Figure 1(a), our DocLayLLM demonstrates superior document understanding capabilities, delivering accurate answers in examples from InfoVQA and VisualMRC. Furthermore, in the DeepForm example, we observe that our design to integrate OCR information helps reduce the occurrence of hallucinated outputs compared to OCR-free methods. Moreover, in the examples from DocVQA and WTQ, DocLayLLM reliably exhibits robust table comprehension abilities. These results collectively highlight the effectiveness of our design in incorporating OCR information.
Furthermore, we also present the results of whether Table-Structure-Aware CoT is used during the pre-training stage. As shown in Figure 1(b), models incorporating CoT demonstrate a more comprehensive understanding of table structures, leading to more accurate and thorough answers in table-related downstream document understanding tasks. This validates the effectiveness of our proposed CoT pre-training approach.
Additionally, we visualized the outputs with and without the use of CoT Annealing. As shown in the visualization of VisualMRC in Figure 1(c), DocLayLLM employing CoT Annealing tends to provide more straightforward and accurate answers. This is particularly evident in yes-or-no questions, where the model without CoT Annealing often fails to directly respond with a clear "yes" or "no" but repeats the sentence in the document where the answer is located. In contrast, the model using CoT Annealing typically provides a direct answer first, followed by an explanation. These observations indicate that CoT Annealing enables the model to answer questions more directly, thereby enhancing its performance.
@@ C More Qualitative Examples
(a) Qualitative comparisons with DocOwl 1.5 [17] across various benchmarks. The document-oriented VQA tasks include InfoVQA [47], VisualMRC [59], and DocVQA [46]; the KIE task includes DeepForm [58]; and the Table Understanding task includes WTQ [51].
(b) Qualitative comparisons between the use and absence of CoT Pre-training. w/o Pre-training indicates the absence of CoT at the pre-training stage, while w/ CoT Pre-training denotes its application. ?" represents that the answer is ambiguous.
(c) Qualitative comparisons between the use and absence of CoT Annealing. w/o CoT Annealing indicates the absence of CoT Annealing, while w/ CoT Annealing denotes its application. ?" represents that the answer is ambiguous.
Figure A: Further qualitative comparisons of DocLayLLM against the SOTA OCR-free method and under various settings.
D Input Length Analysis
In our ablation study, we evaluated the performance of different methods for incorporating OCR information. This section further examines the input length of OCR information under various approaches. The analysis was conducted using Llama3 [42] as the base model, with its tokenizer applied for tokenization. Table C presents a comparison of the average input length of OCR information across several benchmarks under two configurations: (I) encoding OCR bounding box coordinates as plain text, following the approach of ICL-D3IE [14], and (II) encoding OCR bounding box coordinates using a layout embedder $LE$.
The results clearly show that encoding with $LE$ significantly reduces the input length, thereby enhancing efficiency during both training and inference. These findings underscore the efficiency of our proposed DocLayLLM.
Input Method Document-oriented VQA KIE
DocVQA VisualMRC DeepForm KLC
(I) 1571.80 6269.17 4952.87 457.58
(II) 455.17 2095.35 1198.97 125.40
Table C: The average input length of OCR information across various benchmarks, comparing different ways to input OCR bounding box coordinates.
E OCR Result Impacts
Since DocLayLLM requires OCR result input, we explored the impact of OCR quality on the performance of DocLayLLM. In the results presented in the main text, we used the official OCR results when evaluating on the DocVQA benchmark. To assess the model’s applicability in real-world scenarios, we employed a commercial OCR engine  [^2] https://www.textin.com/ to process DocVQA and used the recognized text for further testing DocLayLLM’s performance. The results in Table D suggest that the reported results in the main text have not fully reflected the potential of DocLayLLM. The model could achieve even better performance with real-world OCR results.
Furthermore, as illustrated in Figure B, we observed that when OCR errors occur, DocLayLLM has the capability to correct these errors and produce the final correct answer. This further substantiates the robustness of DocLayLLM in real-world scenarios.
Figure B: Illustration of DocLayLLM’s OCR error correction capability.
[…]
Official OCR 86.52
Commercial Engine 87.52
Table D: Performances of DocLayLLM on DocVQA benchmark with different OCR results.